OpenStudy (anonymous):

find the laplace transforms when f(t)=e^(6t) if 05 Can anyone help?

OpenStudy (amistre64):

\[\large\left.L\{e^{6t}\}\right|_{0}^{5}+\left.L\{1\}\right|_{5}^{inf}\]

OpenStudy (amistre64):

\[e^{6t}\to\ e^{(6-s)t} =\frac{e^{(6-s)5}}{6-s}-\frac{e^{(6-s)0}}{6-s}\]\[1\to\ e^{-st} =\frac{e^{-sinf}}{-s}-\frac{e^{-s5}}{-s}\]

OpenStudy (anonymous):

isnt e^-sinf be equal to zero then?

OpenStudy (amistre64):

yes, unless theyve altered reality when i wasnt looking :)

OpenStudy (anonymous):

so the answer would be e^((6-s)5)/(6-s)-1/(6-s)+e^(-s5)/s would that be correct?

OpenStudy (amistre64):

that last bit should be a - and you might wana simplify for good measure; but yes

OpenStudy (amistre64):

i see, you took the - from teh s .. ok

OpenStudy (anonymous):

well you are deviding by -s so itd be positive. Thanks for the help though! :) It is greatly appreciated :)

OpenStudy (amistre64):

yw :)